# On-board Satellite Image Classification for Earth Observation: A Comparative Study of Pre-trained Vision-Transformer Models

### v1

@@ VI Conclusions
In conclusion, the EfficientViT-M2 model is optimal for implementing IC in On-Air EO processing missions. Its superior robustness to noise, demonstrated across varying levels of Gaussian noise and motion blur, ensures reliable performance in the challenging conditions often encountered in space. Additionally, EfficientViT-M2’s low computational complexity (203.53 MFLOPs) and compact size (38.19 MB) make it highly suitable for resource-constrained environments, offering significant efficiency gains over pioneering models used in previous satellite missions like $\Phi$-Sat-1 and $\Phi$-Sat-2. The model’s ability to maintain high accuracy and classification performance with minimal computational overhead further cements its practicality for real-world deployment, where consistent and accurate image processing is critical. As satellite EO demands increasingly sophisticated yet efficient algorithms, EfficientViT-M2 stands out as a leading solution, combining advanced ML capabilities with the feasibility of successful on-board implementation.
Limitations and Future Works:
One limitation of this study is the challenge of adapting deep neural networks to the dynamic nature of wireless communication systems. Frequent changes in the environment and data distribution can necessitate on-device retraining, adding complexity. Future work should address this by focusing on managing lossy and dynamic transmission errors through practical communication system modeling.
Another limitation is the study’s exclusive focus on IC, without exploring the potential of multitask learning, as exemplified by the Multi-Tasks Pre-trained Model (MTP) [45]. Although the EfficientViT achieves slightly lower accuracy compared to the MTP model’s 99.30%, the MTP model’s incorporation of both CNNs and Transformers results in over 300 million parameters and significantly higher computational complexity, limiting its practicality in resource-constrained environments like on-board satellite systems. Future research should investigate the application of EfficientViT in multitasking scenarios to balance performance with computational efficiency.
Finally, the study is limited by its focus on single-modality data, potentially restricting its ability to leverage the expanding wealth of multimodal EO data. Future research should explore the potential of transformer-based approaches within a multimodal DL framework to utilize better and integrate diverse types of EO data, as demonstrated by [46, 47].
Acknowledgment
This work was funded by the Luxembourg National Research Fund (FNR), with granted SENTRY project corresponding to grant reference C23/IS/18073708/SENTRY.

### v2

@@ VI Conclusions
In this paper, we study and conduct extensive experiments comparing traditional CNN-based, ResNet-based, and Transformer-based models for IC. Our analysis delves into various pre-trained ViT architectures to identify the optimal balance between model complexity, computational efficiency, and performance. A key focus of our evaluation is assessing the robustness of these models under noisy data conditions during inference, which closely mirrors real-world scenarios encountered in satellite EO systems.
In conclusion, the EfficientViT-M2 model is optimal for implementing IC in onboard satellite EO missions. Its superior robustness to noise, demonstrated across varying levels of Gaussian noise and motion blur, ensures reliable performance in the challenging conditions often encountered in space. Additionally, EfficientViT-M2’s low computational complexity (203.53 MFLOPs) and compact size (38.19 MB) make it highly suitable for resource-constrained environments, offering significant efficiency gains over pioneering models used in previous satellite missions like $\Phi$-Sat-1 and $\Phi$-Sat-2. The model’s ability to maintain high accuracy and classification performance with minimal computational overhead further cements its practicality for real-world deployment, where consistent and accurate image processing is critical. As satellite EO demands increasingly sophisticated yet efficient algorithms, EfficientViT-M2 stands out as a leading solution, combining advanced ML capabilities with the feasibility of successful on-board implementation.
One limitation of this study is the challenge of adapting deep NNs to the dynamic nature of wireless communication systems. Frequent changes in the environment and data distribution can necessitate on-device retraining, adding complexity. Future work should address this by focusing on managing lossy and dynamic transmission errors through practical communication system modeling.
Another limitation is the study’s exclusive focus on IC, without exploring the potential of multitask learning, as exemplified by the Multi-Tasks Pre-trained Model (MTP) [52]. Although the EfficientViT achieves slightly lower accuracy compared to the MTP model’s 99.30%, the MTP model’s incorporation of both CNNs and Transformers results in over 300 million parameters and significantly higher computational complexity, limiting its practicality in resource-constrained environments like OSS. Future research should investigate the EfficientViT application in multitasking scenarios for computational efficiency balancing.
Finally, the study is limited by its focus on single-modality data, potentially restricting its ability to leverage the expanding wealth of multimodal EO data. Future research should explore the potential of transformer-based approaches within a multimodal DL framework to utilize better and integrate diverse types of EO data [53, 54].